A rectangle has a perimeter of 40 cm and a length of x cm.
A) show that the area A=(20x-x^2) cm^2
B) if A = 96 cm^2, find the length

Answer by Boreal(15235)

1/2 perimeter= L+W
20=x+W, letting x=length.
W=20-x
A=L*W=x(20-x)
Let it equal 96
20x-x^2=96
x^2-20x+96=0
(x-12)(x-8)=0
x=12, 8 (cm)
The length is considered longer, so it is 12 cm. The width is 8 cm, which gives the right P and A.